# Asymptotic behaviors and dynamics of degenerate and mixed solitons for the coupled Hirota system with strong coherent coupling effects

### v1

@@ 3. Asymptotic behaviors of the degenerate solitons
$\varOmega\left(\Gamma_{d},\Gamma_{j}\right)=-\varOmega\left(\Gamma_{j},\Gamma_{d}\right)^{\dagger}$
$=Y_{d}^{\dagger}Y_{j}\lim_{\lambda_{d}^{*}\to\lambda_{j}}\frac{e^{{\rm i}\left(\eta_{d}^{*}-\eta_{j}\right)}-e^{-{\rm i}\left(\eta_{d}^{*}-\eta_{j}\right)}}{\lambda_{d}^{*}-\lambda_{j}}-M_{j}^{\dagger}$
$=Y_{d}^{\dagger}Y_{j}\left[\frac{2{\rm i}\eta_{j,I}}{\lambda_{j,I}}-8\lambda_{j,I}\left(1+2{\rm i}\lambda_{j,I}\varepsilon+6\lambda_{j,R}\varepsilon\right)t\right]-M_{j}^{\dagger},$
@@
$q_{1}[2]=-2{\rm i}e^{-2{\rm i}\eta_{1,R}}\left(e^{\theta},0,-l_{13}^{*},-l_{14}^{*}\right)\varOmega[2]^{-1}\left(l_{13}e^{-\theta},-l_{14}e^{-\theta},1,0\right)^{\dagger},$ (17a)
$q_{2}[2]=-2{\rm i}e^{-2{\rm i}\eta_{1,R}}\left(e^{\theta},0,-l_{13}^{*},-l_{14}^{*}\right)\varOmega[2]^{-1}\left(l_{14}e^{-\theta},l_{13}e^{-\theta},0,1\right)^{\dagger},$ (17b)
@@
$q_{1}^{\pm}\to\left\{\begin{array}[]{cc}0,&l_{13}^{2}+l_{14}^{2}\not=0,\\ {\rm{i}}\lambda_{1,I}e^{-2{\rm i}\eta_{1,R}}\frac{m_{11}-\frac{l_{13}}{l_{14}}m_{12}}{\left|m_{11}-\frac{l_{13}}{l_{14}}m_{12}\right|}{\rm sech}\left(\theta-\frac{\rho}{2}\right),&l_{13}^{2}+l_{14}^{2}=0,\end{array}\right.$ (20c)
$q_{2}^{\pm}=-\frac{l_{13}}{l_{14}}q_{1}^{\pm},$ (20d)
where the signs “$\pm$” in the superscript correspond to asymptotic limits of the soliton when $t\to\pm\infty$, and $e^{\frac{\rho}{2}}=2\lambda_{1,I}\left|m_{11}-\frac{l_{13}}{l_{14}}m_{12}\right|$. From Asymptotic Expressions (20), we find that the asymptotic limits of Solutions (17) are the sech-type functions under the conditions $l_{13}^{2}+l_{14}^{2}=0$, which implies that Solutions (17) can generate the bell-shaped soliton branch the straight line $\theta=\mathcal{O}(1)$ as $|t|\to\infty$, as shown in Figs. 1.
[Uncaptioned image][Uncaptioned image]
(a)                                                          (b)
@@
$te^{\pm\theta}\sim\mathcal{O}(1).$ (21)
To analyze the asymptotic behaviors of solitons along with Banalance (21), we rewrite Solutions (17) in the form of
$q_{s}=2{\rm i}e^{-2{\rm i}\eta_{1,R}}\frac{{\rm det}\left(F_{s}\right)}{{\rm det}\left(\Omega[2]\right)}=2{\rm i}e^{-2{\rm i}\eta_{1,R}}\frac{{\rm det\left(\begin{array}[]{cc}\Omega[2]&B^{[s]\dagger}\\ A&0\end{array}\right)}}{{\rm det}\left(\Omega[2]\right)},\quad(s=1,2),$ (24)
where $A=\left(e^{\theta},0,-l_{13}^{*},-l_{14}^{*}\right)$, the superscript $[s]$ represents the $s$th row of a matrix, and $B=\left(Y_{1}e^{-\theta},{\bf I}_{2\times 2}\right)$.
Based on Banalance (21) and Solutions (24), we can obtain two distinct asymptotic behaviors of solitons:
@@
$\left(\begin{array}[]{c}q_{1}\\ q_{2}\end{array}\right)\to\left(\begin{array}[]{c}q_{1}\\ q_{2}\end{array}\right)^{(1)}=\frac{2{\rm i}e^{-2{\rm i}\eta_{1,R}}}{e^{4\theta}{\rm det}\left(\tilde{\Omega}[2]^{(1)}\right)}\left(\begin{array}[]{c}e^{4\theta}{\rm det}\left(\tilde{F}_{1}^{(1)}\right)\\ e^{4\theta}{\rm det}\left(\tilde{F}_{2}^{(1)}\right)\end{array}\right)$ (31)
@@
$\tilde{\Omega}[2]^{(1)}=\left(\begin{array}[]{cc}-\frac{{\rm i}}{2\lambda_{1,I}}{\bf I}_{2\times 2}&8\lambda_{1,I}\chi^{*}Y_{1}^{\dagger}\frac{t}{e^{\theta}}\\ -8\lambda_{1,I}\chi Y_{1}\frac{t}{e^{\theta}}&\frac{{\rm i}}{2\lambda_{1,I}}{\bf I}_{2\times 2}\end{array}\right),$
$\tilde{F}_{s}^{(1)}=\left(\begin{array}[]{ccc}-\frac{{\rm i}}{2\lambda_{1,I}}{\bf I}_{2\times 2}&8\lambda_{1,I}\chi^{*}Y_{1}^{\dagger}\frac{t}{e^{\theta}}&{\bf 0}_{2\times 1}\\ -8\lambda_{1,I}\chi Y_{1}\frac{t}{e^{\theta}}&\frac{{\rm i}}{2\lambda_{1,I}}{\bf I}_{2\times 2}&\left({\bf I}_{2\times 2}^{[s]}\right)^{\dagger}\\ {\bf I}_{2\times 2}^{[1]}&{\bf 0}_{1\times 2}&0\end{array}\right),$
$\varphi_{1}^{(1)}=2{\rm i}\lambda_{1,I}e^{-2{\rm i}\eta_{1,R}}\frac{\left(l_{13}-{\rm i}l_{14}\right)^{*}\chi^{*}}{\left|\left(l_{13}-{\rm i}l_{14}\right)\chi\right|}{\rm sech}\left(\xi+\mu_{1}\right),$
$\varphi_{2}^{(1)}=2{\rm i}\lambda_{1,I}e^{-2{\rm i}\eta_{1,R}}\frac{\left(l_{13}+{\rm i}l_{14}\right)^{*}\chi^{*}}{\left|\left(l_{13}+{\rm i}l_{14}\right)\chi\right|}{\rm sech}\left(\xi+\nu_{1}\right),$
where ${\bf 0}_{2\times 1}$ is a $2\times 1$ zero matrix, ${\bf 0}_{1\times 2}$ denotes a $1\times 2$ zero matrix, $e^{\xi}=\frac{8\lambda_{1,I}t}{e^{\theta}}$, $e^{\mu_{1}}=2\lambda_{1,I}\left|\left(l_{13}-{\rm i}l_{14}\right)\chi\right|$ and $e^{\nu_{1}}=2\lambda_{1,I}\left|\left(l_{13}+{\rm i}l_{14}\right)\chi\right|$. Obviously, Asymptotic Expression (36) can be viewed as the linear superposition of two sech-type functions, which implies that Asymptotic Expression (36) can describe the superposition of two bell-shaped solitons. The soliton $\left(\begin{array}[]{c}q_{1}^{(1)}\\ q_{2}^{(1)}\end{array}\right)$ propagates at the velocity
@@
$\left(\begin{array}[]{c}q_{1}\\ q_{2}\end{array}\right)\to\left(\begin{array}[]{c}q_{1}\\ q_{2}\end{array}\right)^{(2)}=\frac{2{\rm i}e^{-2{\rm i}\eta_{1,R}}}{e^{-4\theta}{\rm det}\left(\tilde{\Omega}[2]^{(2)}\right)}\left(\begin{array}[]{c}e^{-4\theta}{\rm det}\left(\tilde{F}_{1}^{(2)}\right)\\ e^{-4\theta}{\rm det}\left(\tilde{F}_{2}^{(2)}\right)\end{array}\right)$ (46)
@@
$\tilde{\Omega}[2]^{(2)}=\left(\begin{array}[]{cc}-\frac{{\rm i}}{2\lambda_{1,I}}Y_{1}^{\dagger}Y_{1}&8\lambda_{1,I}\chi^{*}Y_{1}^{\dagger}te^{\theta}\\ -8\lambda_{1,I}\chi Y_{1}te^{\theta}&\frac{{\rm i}}{2\lambda_{1,I}}Y_{1}Y_{1}^{\dagger}\end{array}\right),$
$\tilde{F}_{s}^{(2)}=\left(\begin{array}[]{ccc}-\frac{{\rm i}}{2\lambda_{1,I}}Y_{1}^{\dagger}Y_{1}&8\lambda_{1,I}\chi^{*}Y_{1}^{\dagger}te^{\theta}&\left(Y_{1}^{[s]}\right)^{\dagger}\\ -8\lambda_{1,I}\chi Y_{1}te^{\theta}&\frac{{\rm i}}{2\lambda_{1,I}}Y_{1}Y_{1}^{\dagger}&{\bf 0}_{2\times 1}\\ {\bf 0}_{1\times 2}&-\left(Y_{1}^{\dagger}\right)^{[1]}&0\end{array}\right),$
$\varphi_{1}^{(2)}=2{\rm i}\lambda_{1,I}e^{-2{\rm i}\eta_{1,R}}\frac{\left(l_{13}-{\rm i}l_{14}\right)^{*}\chi}{\left|\left(l_{13}-{\rm i}l_{14}\right)\chi\right|}{\rm sech}\left(\zeta+\mu_{2}\right),$
$\varphi_{2}^{(2)}=2{\rm i}\lambda_{1,I}e^{-2{\rm i}\eta_{1,R}}\frac{\left(l_{13}+{\rm i}l_{14}\right)^{*}\chi}{\left|\left(l_{13}+{\rm i}l_{14}\right)\chi\right|}{\rm sech}\left(\zeta+\nu_{2}\right),$
@@
$S^{1-}=\left(\begin{array}[]{c}q_{1}^{(1)-}\\ q_{2}^{(1)-}\end{array}\right),\quad S^{1+}=\left(\begin{array}[]{c}q_{1}^{(2)+}\\ q_{2}^{(2)+}\end{array}\right),$ (55e)
$S^{2-}=\left(\begin{array}[]{c}q_{1}^{(2)-}\\ q_{2}^{(2)-}\end{array}\right),\quad S^{2+}=\left(\begin{array}[]{c}q_{1}^{(1)+}\\ q_{2}^{(1)+}\end{array}\right),$ (55j)
@@
Analyzing Asymptotic Expressions (55), we see that $\left|q_{1}\right|$ and $\left|q_{2}\right|$ are not proportional in the two vector asymptotic solitons $S^{1}$ and $S^{2}$, which reflects that $S^{1}$ and $S^{2}$ admit the distinct intensity profiles in the $q_{1}$ and $q_{2}$ components. In addition, the two soliton branches of $S^{1}$ and $S^{2}$ defined by Asymptotic Expressions (55) can be split up with $m_{11}$ increasing, as seen in Figs. 2.
@@
(a1)                                                         (b1)
(a2)                                                         (b2)
Based on expressions of $\mu_{1}$, $\mu_{2}$, $\nu_{1}$ and $\nu_{2}$, we find that $\left|\mu_{1}-\nu_{1}\right|=\left|\mu_{2}-\nu_{2}\right|$, which indicates that the profiles of $S^{1}$ and $S^{2}$ are the same and remain invariant both before and after the interactions. For instance, Figs. 3 show the interaction between two-soliton branches: Figs. 3(a1-a2) depict the interaction between two single-hump solitons in the $q_{1}$ component, while Figs. 3(b1-b2) display that two double-hump solitons interact with each other in the $q_{2}$ component. Each of $q_{1}$ and $q_{2}$ components may exhibit peak profiles in their interaction regions.
@@
$S^{\rm line-}=S^{\rm line+}=\left(\begin{array}[]{c}q_{1}^{\rm line}\\ q_{2}^{\rm line}\end{array}\right),$ (56c)
$S^{1-}=\left(\begin{array}[]{c}q_{1}^{(1)-}\\ q_{2}^{(1)-}\end{array}\right),\quad S^{1+}=\left(\begin{array}[]{c}q_{1}^{(2)+}\\ q_{2}^{(2)+}\end{array}\right),$ (56h)
$S^{2-}=\left(\begin{array}[]{c}q_{1}^{(2)-}\\ q_{2}^{(2)-}\end{array}\right),\quad S^{2+}=\left(\begin{array}[]{c}q_{1}^{(1)+}\\ q_{2}^{(1)+}\end{array}\right),$ (56m)
@@
$q_{1}^{(1)}={\rm i}\lambda_{1,I}e^{-2{\rm i}\eta_{1,R}}\frac{l_{13}^{*}\chi^{*}}{\left|l_{13}\chi\right|}{\rm sech}\left(\xi+\mu\right),\quad q_{2}^{(1)}=\frac{l_{13}}{l_{14}}q_{1}^{(1)},$
$q_{1}^{(2)}={\rm i}\lambda_{1,I}e^{-2{\rm i}\eta_{1,R}}\frac{l_{13}^{*}\chi}{\left|l_{13}\chi\right|}{\rm sech}\left(\zeta+\nu\right),\quad q_{2}^{(2)}=\frac{l_{13}}{l_{14}}q_{1}^{(2)},$
$q_{1}^{\rm line}={\rm{i}}\lambda_{1,I}e^{-2{\rm i}\eta_{1,R}}\frac{m_{11}-\frac{l_{13}}{l_{14}}m_{12}}{\left|m_{11}-\frac{l_{13}}{l_{14}}m_{12}\right|}{\rm sech}\left(\theta-\frac{\rho}{2}\right),\quad q_{2}^{\rm line}=-\frac{l_{13}}{l_{14}}q_{1}^{\rm line},$
@@
Figs. 4(a1-a2) and 4(b1-b2) display the elastic interaction among the three vector solitons $S^{1}$, $S^{2}$ and $S^{\rm line}$. Obviously, the intensity profiles of $S^{1}$, $S^{2}$ and $S^{\rm line}$ in both the $q_{1}$ and $q_{2}$ components exhibit similarity due to their proportionality. In addition, during the process of the interaction, the profile of the line soliton $S^{\rm line}$ remains unchanged, with no observed phase shifts.
@@
Based on Expression (63), when the values of $t$ and $\lambda_{1,I}$ are fixed, the relative distance between the two asymptotic solitons $S^{1}$ and $S^{2}$ increases logarithmically with the higher-order perturbation parameter $|\varepsilon|$, as shown in Figs. 5.
@@
  [Uncaptioned image]   [Uncaptioned image]
@@
In the three cases discussed above, the asymptotic solitons can be localized along the straight line or certain algebraic curves according to Solutions (24), all of which can be approached by the exact solutions when $|t|$ tends to infinity. In what follows, we will graphically demonstrate the validity of the asymptotic analysis. In Figs. 7-9, we compare the degenerate solitons described by Solutions (24) with the asymptotic solitons given in Asymptotic Expressions (55) and (56) for a large value of $|t|$.
@@
Figs. 7-8 show that all the asymptotic solitons closely match the degenerate solitons when $t=30$, which confirms the accuracy of our asymptotic analysis in describing soliton behaviors.

### v2

@@ 3. Asymptotic behaviors of the degenerate solitons
$\hskip-56.9055pt\varOmega\left(\Gamma_{d},\Gamma_{j}\right)=-\varOmega\left(\Gamma_{j},\Gamma_{d}\right)^{\dagger}$
$\hskip-4.26773pt=Y_{d}^{\dagger}Y_{j}\lim_{\lambda_{d}^{*}\to\lambda_{j}}\frac{e^{{\rm i}\left(\eta_{d}^{*}-\eta_{j}\right)}-e^{-{\rm i}\left(\eta_{d}^{*}-\eta_{j}\right)}}{\lambda_{d}^{*}-\lambda_{j}}-M_{j}^{\dagger}$
$\hskip-4.26773pt=Y_{d}^{\dagger}Y_{j}\left[\frac{2{\rm i}\eta_{j,I}}{\lambda_{j,I}}-8\lambda_{j,I}\left(1+2{\rm i}\lambda_{j,I}\varepsilon+6\lambda_{j,R}\varepsilon\right)t\right]-M_{j}^{\dagger},$
@@
$\hskip-28.45274ptq_{1}[2]=-2{\rm i}e^{-2{\rm i}\eta_{1,R}}\left(e^{\theta},0,-l_{13}^{*},-l_{14}^{*}\right)\varOmega[2]^{-1}\left(l_{13}e^{-\theta},-l_{14}e^{-\theta},1,0\right)^{\dagger},$ (17a)
$\hskip-28.45274ptq_{2}[2]=-2{\rm i}e^{-2{\rm i}\eta_{1,R}}\left(e^{\theta},0,-l_{13}^{*},-l_{14}^{*}\right)\varOmega[2]^{-1}\left(l_{14}e^{-\theta},l_{13}e^{-\theta},0,1\right)^{\dagger},$ (17b)
@@
$\hskip-42.67912ptq_{1}^{\pm}\to\left\{\begin{array}[]{cc}0,&l_{13}^{2}+l_{14}^{2}\not=0,\\ {\rm{i}}\lambda_{1,I}e^{-2{\rm i}\eta_{1,R}}\frac{m_{11}-\frac{l_{13}}{l_{14}}m_{12}}{\left|m_{11}-\frac{l_{13}}{l_{14}}m_{12}\right|}{\rm sech}\left(\theta-\frac{\rho}{2}\right),&l_{13}^{2}+l_{14}^{2}=0,\end{array}\right.$ (20c)
$\hskip-42.67912ptq_{2}^{\pm}=-\frac{l_{13}}{l_{14}}q_{1}^{\pm},$ (20d)
where the signs “$\pm$” in the superscript correspond to asymptotic limits of the soliton when $t\to\pm\infty$, and $e^{\frac{\rho}{2}}=2\lambda_{1,I}\left|m_{11}-\frac{l_{13}}{l_{14}}m_{12}\right|$. From Asymptotic Expressions (20), we find that the asymptotic limits of Solutions (17) are the sech-type functions under the conditions $l_{13}^{2}+l_{14}^{2}=0$, which implies that Solutions (17) can generate the bell-shaped soliton branch the straight line $\theta=\mathcal{O}(1)$ as $|t|\to\infty$.
@@
$\hskip-28.45274ptte^{\pm\theta}\sim\mathcal{O}(1).$ (21)
To analyze the asymptotic behaviors of solitons along with Expression (21), we rewrite Solutions (17) in the form of
$\hskip-36.98866ptq_{s}=2{\rm i}e^{-2{\rm i}\eta_{1,R}}\frac{{\rm det}\left(F_{s}\right)}{{\rm det}\left(\Omega[2]\right)}=2{\rm i}e^{-2{\rm i}\eta_{1,R}}\frac{{\rm det\left(\begin{array}[]{cc}\Omega[2]&B^{[s]\dagger}\\ A&0\end{array}\right)}}{{\rm det}\left(\Omega[2]\right)},\quad(s=1,2),$ (24)
[…]
Based on Expression (21) and Solutions (24), we can obtain two distinct asymptotic behaviors of solitons:
@@
$\hskip-42.67912pt\left(\begin{array}[]{c}q_{1}\\ q_{2}\end{array}\right)\to\left(\begin{array}[]{c}q_{1}\\ q_{2}\end{array}\right)^{(1)}=\frac{2{\rm i}e^{-2{\rm i}\eta_{1,R}}}{e^{4\theta}{\rm det}\left(\tilde{\Omega}[2]^{(1)}\right)}\left(\begin{array}[]{c}e^{4\theta}{\rm det}\left(\tilde{F}_{1}^{(1)}\right)\\ e^{4\theta}{\rm det}\left(\tilde{F}_{2}^{(1)}\right)\end{array}\right)$ (31)
@@
$\hskip-2.84544pt\tilde{\Omega}[2]^{(1)}=\left(\begin{array}[]{cc}-\frac{{\rm i}}{2\lambda_{1,I}}{\bf I}_{2\times 2}&8\lambda_{1,I}\chi^{*}Y_{1}^{\dagger}\frac{t}{e^{\theta}}\\ -8\lambda_{1,I}\chi Y_{1}\frac{t}{e^{\theta}}&\frac{{\rm i}}{2\lambda_{1,I}}{\bf I}_{2\times 2}\end{array}\right),$
$\hskip-2.84544pt\tilde{F}_{s}^{(1)}=\left(\begin{array}[]{ccc}-\frac{{\rm i}}{2\lambda_{1,I}}{\bf I}_{2\times 2}&8\lambda_{1,I}\chi^{*}Y_{1}^{\dagger}\frac{t}{e^{\theta}}&{\bf 0}_{2\times 1}\\ -8\lambda_{1,I}\chi Y_{1}\frac{t}{e^{\theta}}&\frac{{\rm i}}{2\lambda_{1,I}}{\bf I}_{2\times 2}&\left({\bf I}_{2\times 2}^{[s]}\right)^{\dagger}\\ {\bf I}_{2\times 2}^{[1]}&{\bf 0}_{1\times 2}&0\end{array}\right),$
$\hskip-2.84544pt\varphi_{1}^{(1)}=2{\rm i}\lambda_{1,I}e^{-2{\rm i}\eta_{1,R}}\frac{\left(l_{13}-{\rm i}l_{14}\right)^{*}\chi^{*}}{\left|\left(l_{13}-{\rm i}l_{14}\right)\chi\right|}{\rm sech}\left(\xi+\mu_{1}\right),$
$\hskip-2.84544pt\varphi_{2}^{(1)}=2{\rm i}\lambda_{1,I}e^{-2{\rm i}\eta_{1,R}}\frac{\left(l_{13}+{\rm i}l_{14}\right)^{*}\chi^{*}}{\left|\left(l_{13}+{\rm i}l_{14}\right)\chi\right|}{\rm sech}\left(\xi+\nu_{1}\right),$
where ${\bf 0}_{2\times 1}$ is a 2×12$\times$ 12 × 1 zero matrix, ${\bf 0}_{1\times 2}$ denotes a 1×21$\times$ 21 × 2 zero matrix, $e^{\xi}=\frac{8\lambda_{1,I}t}{e^{\theta}}$, $e^{\mu_{1}}=2\lambda_{1,I}\left|\left(l_{13}-{\rm i}l_{14}\right)\chi\right|$ and $e^{\nu_{1}}=2\lambda_{1,I}\left|\left(l_{13}+{\rm i}l_{14}\right)\chi\right|$. Obviously, Asymptotic Expression (36) can be viewed as the linear superposition of two sech-type functions, which implies that Asymptotic Expression (36) can describe the superposition of two bell-shaped solitons. The soliton $\left(\begin{array}[]{c}q_{1}^{(1)}\\ q_{2}^{(1)}\end{array}\right)$ propagates at the velocity
@@
$\hskip-42.67912pt\left(\begin{array}[]{c}q_{1}\\ q_{2}\end{array}\right)\to\left(\begin{array}[]{c}q_{1}\\ q_{2}\end{array}\right)^{(2)}=\frac{2{\rm i}e^{-2{\rm i}\eta_{1,R}}}{e^{-4\theta}{\rm det}\left(\tilde{\Omega}[2]^{(2)}\right)}\left(\begin{array}[]{c}e^{-4\theta}{\rm det}\left(\tilde{F}_{1}^{(2)}\right)\\ e^{-4\theta}{\rm det}\left(\tilde{F}_{2}^{(2)}\right)\end{array}\right)$ (46)
@@
$\hskip-2.84544pt\tilde{\Omega}[2]^{(2)}=\left(\begin{array}[]{cc}-\frac{{\rm i}}{2\lambda_{1,I}}Y_{1}^{\dagger}Y_{1}&8\lambda_{1,I}\chi^{*}Y_{1}^{\dagger}te^{\theta}\\ -8\lambda_{1,I}\chi Y_{1}te^{\theta}&\frac{{\rm i}}{2\lambda_{1,I}}Y_{1}Y_{1}^{\dagger}\end{array}\right),$
$\hskip-2.84544pt\tilde{F}_{s}^{(2)}=\left(\begin{array}[]{ccc}-\frac{{\rm i}}{2\lambda_{1,I}}Y_{1}^{\dagger}Y_{1}&8\lambda_{1,I}\chi^{*}Y_{1}^{\dagger}te^{\theta}&\left(Y_{1}^{[s]}\right)^{\dagger}\\ -8\lambda_{1,I}\chi Y_{1}te^{\theta}&\frac{{\rm i}}{2\lambda_{1,I}}Y_{1}Y_{1}^{\dagger}&{\bf 0}_{2\times 1}\\ {\bf 0}_{1\times 2}&-\left(Y_{1}^{\dagger}\right)^{[1]}&0\end{array}\right),$
$\hskip-2.84544pt\varphi_{1}^{(2)}=2{\rm i}\lambda_{1,I}e^{-2{\rm i}\eta_{1,R}}\frac{\left(l_{13}-{\rm i}l_{14}\right)^{*}\chi}{\left|\left(l_{13}-{\rm i}l_{14}\right)\chi\right|}{\rm sech}\left(\zeta+\mu_{2}\right),$
$\hskip-2.84544pt\varphi_{2}^{(2)}=2{\rm i}\lambda_{1,I}e^{-2{\rm i}\eta_{1,R}}\frac{\left(l_{13}+{\rm i}l_{14}\right)^{*}\chi}{\left|\left(l_{13}+{\rm i}l_{14}\right)\chi\right|}{\rm sech}\left(\zeta+\nu_{2}\right),$
@@
$\hskip-28.45274ptS^{1-}=\left(\begin{array}[]{c}q_{1}^{(1)-}\\ q_{2}^{(1)-}\end{array}\right),\quad S^{1+}=\left(\begin{array}[]{c}q_{1}^{(2)+}\\ q_{2}^{(2)+}\end{array}\right),$ (55e)
$\hskip-28.45274ptS^{2-}=\left(\begin{array}[]{c}q_{1}^{(2)-}\\ q_{2}^{(2)-}\end{array}\right),\quad S^{2+}=\left(\begin{array}[]{c}q_{1}^{(1)+}\\ q_{2}^{(1)+}\end{array}\right),$ (55j)
@@
Analyzing Asymptotic Expressions (55), we see that $\left|q_{1}\right|$ and $\left|q_{2}\right|$ are not proportional in the two vector asymptotic solitons $S^{1}$ and $S^{2}$, which reflects that $S^{1}$ and $S^{2}$ admit the distinct intensity profiles in the $q_{1}$ and $q_{2}$ components as seen in Figs. 1.
@@
(a)                    (b)
Based on expressions of $\mu_{1}$, $\mu_{2}$, $\nu_{1}$ and $\nu_{2}$, we find that $\left|\mu_{1}-\nu_{1}\right|=\left|\mu_{2}-\nu_{2}\right|$, which indicates that the profiles of $S^{1}$ and $S^{2}$ are the same and remain invariant both before and after the interactions. For instance, Figs. 2 show the interaction between two-soliton branches: Figs. 2(a1-a2) depict the interaction between two single-hump solitons in the $q_{1}$ component, while Figs. 2(b1-b2) display that two double-hump solitons interact with each other in the $q_{2}$ component. Both $q_{1}$ and $q_{2}$ components exhibit peak profiles in their interaction regions.
@@
$\hskip-28.45274ptS^{\rm line-}=S^{\rm line+}=\left(\begin{array}[]{c}q_{1}^{\rm line}\\ q_{2}^{\rm line}\end{array}\right),$ (56c)
$\hskip-28.45274ptS^{1-}=\left(\begin{array}[]{c}q_{1}^{(1)-}\\ q_{2}^{(1)-}\end{array}\right),\quad S^{1+}=\left(\begin{array}[]{c}q_{1}^{(2)+}\\ q_{2}^{(2)+}\end{array}\right),$ (56h)
$\hskip-28.45274ptS^{2-}=\left(\begin{array}[]{c}q_{1}^{(2)-}\\ q_{2}^{(2)-}\end{array}\right),\quad S^{2+}=\left(\begin{array}[]{c}q_{1}^{(1)+}\\ q_{2}^{(1)+}\end{array}\right),$ (56m)
@@
$\hskip-28.45274ptq_{1}^{(1)}={\rm i}\lambda_{1,I}e^{-2{\rm i}\eta_{1,R}}\frac{l_{13}^{*}\chi^{*}}{\left|l_{13}\chi\right|}{\rm sech}\left(\xi+\mu\right),\quad q_{2}^{(1)}=\frac{l_{13}}{l_{14}}q_{1}^{(1)},$
$\hskip-28.45274ptq_{1}^{(2)}={\rm i}\lambda_{1,I}e^{-2{\rm i}\eta_{1,R}}\frac{l_{13}^{*}\chi}{\left|l_{13}\chi\right|}{\rm sech}\left(\zeta+\nu\right),\quad q_{2}^{(2)}=\frac{l_{13}}{l_{14}}q_{1}^{(2)},$
$\hskip-28.45274ptq_{1}^{\rm line}={\rm{i}}\lambda_{1,I}e^{-2{\rm i}\eta_{1,R}}\frac{m_{11}-\frac{l_{13}}{l_{14}}m_{12}}{\left|m_{11}-\frac{l_{13}}{l_{14}}m_{12}\right|}{\rm sech}\left(\theta-\frac{\rho}{2}\right),\quad q_{2}^{\rm line}=-\frac{l_{13}}{l_{14}}q_{1}^{\rm line},$
@@
Figs. 3(a1-a2) and 3(b1-b2) display the elastic interaction among the three vector solitons $S^{1}$, $S^{2}$ and $S^{\rm line}$. Obviously, the intensity profiles of $S^{1}$, $S^{2}$ and $S^{\rm line}$ in both the $q_{1}$ and $q_{2}$ components exhibit similarity due to their proportionality. In addition, during the process of the interaction, the profile of the line soliton $S^{\rm line}$ remains unchanged, with no observed phase shifts.
@@
Based on Expression (63), when the values of $t$ and $\lambda_{1,I}$ are fixed, the relative distance between the two asymptotic solitons $S^{1}$ and $S^{2}$ increases logarithmically with the higher-order perturbation parameter $|\varepsilon|$, as shown in Figs. 4. It is noted that the relative distance between such solitons is always unchanged for the scalar NLS equation and the coupled NLS systems when the eigenvalue $\lambda_{1}$ is fixed.
@@
[Uncaptioned image][Uncaptioned image][Uncaptioned image]
@@
In the three cases discussed above, the asymptotic solitons can be localized along the straight line or certain algebraic curves according to Solutions (24), all of which can be approached by the exact solutions when $|t|$ tends to infinity. In what follows, we will graphically demonstrate the validity of the asymptotic analysis. In Figs. 5-7, we compare the degenerate solitons described by Solutions (24) with the asymptotic solitons given in Asymptotic Expressions (55) and (56) for a large value of $|t|$.
@@
Figs. 6-7 show that all the asymptotic solitons closely match the degenerate solitons when $t=30$, which confirms the accuracy of our asymptotic analysis in describing soliton behaviors.
@@
In the above discussion, to simplify the mathematical analysis and focus on certain physical phenomena such as the coherent effect and robustness of solitons, we make certain specific choices and restrictions on the parameters including $l_{13}$, $l_{14}$, $\lambda_{1}$, and $\varepsilon$ in this paper. These choices enable us to derive the degenerate soliton solutions in a manageable form, but also limit the scope of the solution space. A broader solution space may include other parameter values, which could lead to different types of soliton behaviors or more complex interactions. Our future research will relax these restrictions and explore a broader solution space to reveal new types of solitons and interaction mechanisms, thereby providing a more comprehensive understanding of the physical behavior of the system.